OpenStudy (anonymous):

ln(x^3-1)-ln(x^2+x+1) express in single logarithm with coefficient of 1

OpenStudy (anonymous):

\[\ln \frac{ (x ^{3}-1) }{x ^{2}+x+1 }\]

OpenStudy (anonymous):

now what? do i need to expand the denominator?

OpenStudy (anonymous):

x^-1=(x-1)(x^2+x+1)

OpenStudy (klimenkov):

Better to expand the numerator.

OpenStudy (anonymous):

i gave u the expansion

OpenStudy (anonymous):

why is the expansion of x^3-1 = (x^2+x+1)? I am probably doing it wrong, but dont you pull out x-1 so you would have (x-1)(x^2-1)?

OpenStudy (anonymous):

pull out x-1 then u have x^2+x+1

OpenStudy (anonymous):

wait, if you pull out (x-1) is it (x-1)(x^2+1)?

OpenStudy (anonymous):

\[=\ln (x-1)\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes abselutely